algoritmos_{2}011, type =

Master′sT hesis, title = AlgoritmosdeControlodoMovimentoparaumRobHumanoide, school = UniversityofAveiro, author =

Lage, W illiam, year = 2011,

desenvolvimento_{2}011, type = Master′sT hesis, title = DesenvolvimentodoT roncoeBraçosdeumaP lataformaHumanoideHíbrida, school = UniversityofAveiro, author = Godinho, Ricardo, year = 2011,

virtual_{2}011, title = V irtualRealityandP rogrammingbyDemonstration : T eachingaRobottoGraspaDynamicObjectbytheGeneralizationofHumanDemonstrations, volume = 20, issn = 1054 − 7460, shorttitle = V irtualRealityandP rogrammingbyDemonstration, url = http : //dx.doi.org/10.1162/P RES_{a}_{0}0047, doi = 10.1162/P RES_{a}_{0}0047, number = 3, journal = P resence : T eleoperatorsandV irtualEnvironments, author =

Hamon, LudovicandLucidarme, P hilippeandRichard, EmmanuelleandRichard, P aul, year = 2011, pages = 241 − −253,

hip_{2}007, title = Hip, knee, anklekinematicsandkineticsduringstairascentanddescentinhealthyyoungindividuals, volume = 22, issn = 0268 − 0033, doi = 10.1016/j.clinbiomech.2006.09.010, number = 2, journal = ClinicalBiomechanics, author = P rotopapadaki, AnastasiaandDrechsler, W endyI.andCramp, MaryC.andCoutts, F ionaJ.andScott, OonaM., month = feb, year = 2007, keywords =

Kinematics, Kinetics, Motionanalysis, Stairclimbing, pages = 203 − −210,

modeling_{2}006, address = Berlin, Heidelberg, title = ModelingandSimulationofHumanoidStairClimbing, isbn = 978 − 3 − 540 − 26413 − 2, 978 − 3 − 540 − 26415 − 6, booktitle = ClimbingandW alkingRobots, publisher = SpringerBerlinHeidelberg, author = W ang, L.andZhao, Y.andT okhi, M.O.andGharooni, S.C., editor =

T okhi, M.O.andV irk, G.S.andHossain, M.A., year = 2006, pages =

73 − −80, file = W angetal. − 2006 − ModelingandSimulationofHumanoidStairClimbing.html : /home/pedro/.zotero/zotero/8b71mw6w.default/zotero/storage/B2JT AXF X/W angetal. − 2006 − ModelingandSimulationofHumanoidStairClimbing.html : text/html,

transfer_{2}010, address = Anchorage, Alaska, USA, title = T ransferofskillsbetweenhumanoperatorsthroughhaptictrainingwithrobotcoordination, isbn = 978 − 1 − 4244 − 5038 − 1, url = http : //ieeexplore.ieee.org/xpls/abs_{a}ll.jsp?arnumber = 5509160, doi = 10.1109/ROBOT.2010.5509160, urldate = 2012 − 05 − 01, booktitle = 2010IEEEInternationalConferenceonRoboticsandAutomation(ICRA), publisher = IEEE, author =

P ark, ChungHyukandY oo, JaeW ookandHoward, AyannaM, month = may, year = 2010, pages = 229 − −235,

haptic_{2}002, title = Hapticinterfacesfortheremotecontrolofmobilerobots, volume = 10, issn = 0967 − 0661, doi = 10.1016/S0967 − 0661(02)00153 − 3, number = 11, journal = ControlEngineeringP ractice, author =

Rsch, OttoJandSchilling, KlausandRoth, Hubert, month = nov, year =

2002, keywords =

F orcefeedback, Hapticinterfaces, Mobilerobots, Remotecontrol, Sensorsformobilerobots, pages = 1309 − −1313,

haptic_{2}004, title = Hapticinterfacesanddevices, volume = 24, doi = 10.1108/02602280410515770, number = 1, journal = SensorReview, author = Hayward, V.andAstley, O.R.andCruz − Hernandez, M.andGrant, D.andRobles − De − La − T orre, G., year = 2004, keywords =

Communications, Robotics, Sensors, UserInterfaces, pages = 16 − −29,

hybrid_{2}011, title = Hybridactuationanddistributedcontrolforasmall − sizehumanoidrobot, booktitle = IntroductiontoModernRobotics, publisher = iConceptP ress, author = Santos, V torandMoreira, RuiandSilva, F ilipe, editor = D.ChugoandS.Y okoda, year = 2011, keywords = Humanoidrobot, Humanoidrobotics, Hybridactuation,

development_{2}010, title = DevelopmentofaHybridHumanoidP latformandIncorporationoftheP assiveActuators, booktitle = P roceedingsoftheIEEEInternationalConferenceonRoboticsandBiomimetics, author =

Santos, V torandMoreira, RuiandRibeiro, MiguelandSilva, F ilipe, year = 2010, pages = 679 − −684,

estrutura_{2}009, type = Master′sT hesis, title = EstruturaHíbridadeLocomoçãoparaumRobôHumanoide, school = UniversityofAveiro, author = Sabino, Rmi, year = 2009,

desenvolvimento_{2}010, type = Master′sT hesis, title = DesenvolvimentodosSistemasSensorialeMotorP araUmRobôHumanoide, school = UniversityofAveiro, author = Ribeiro, Miguel, year = 2010,

tele− kinesthetic_{2}012, address =

V ilamoura, P ortugal, title = T ele − KinestheticT eachingofaHumanoidRobotwithHapticDataAcquisition, booktitle = IntelligentRobotsandSystems(IROS) :LearningandInteractioninHapticRobotsW orkshop, author =

Cruz, P edroandSantos, V torandSilva, F ilipe, year = 2012,

phantom_{2}008, address = Boston, USA, title = P HANT oMOMNIUser′sGuide, publisher =

SensAbleT echnologies, year = 2008,